OpenStudy (anonymous):

if a pro basketball player has a vertical leap of about 35inches what is his hangtime use v=48t^2 simplify and use either integer or decimal rounded to nearest 10th

OpenStudy (sriram):

dx/dt=48t^2 integrating put limit of x from 0 to 35 and time from 0 to t x=48t^3/3 x=16t^3 35=16t^3 t=(35/16)^1/3

OpenStudy (sriram):

1.29 approx